# Intuitionistic NF set theory

### v5

@@ 6. USC, SSC, and similarity
[…]
Then we need only consider unit classes $t=\{x\}$, and using the fact that $\{x\}\in USC(b){\ \leftrightarrow\ }t\in b$, and $\{x\}\in USC(a){\ \leftrightarrow\ }t\in a$, (52) follows from 50. That completes the proof of the lemma.
@@
$a\in SSC(b){\ \leftrightarrow\ }SSC(a)\subseteq SSC(b).$
Proof. Left to right: Suppose $a\in SSC(b)$. Then $a\subseteq b$ and
(53) $b$ $=$ $a\cup(b-a)$
Now let $x\in SSC(a)$. We must show $x\in SSC(b)$. Since $x\in SSC(a)$, we have $x\subseteq a$. Since $a\subseteq b$ we have $x\subseteq b$. We have
$x\in SSC(a)$
$a=x\cup(a-x)$   by definition of $SSC(a)$
@@
Right to left: Suppose $SSC(a)\subseteq SSC(b)$. We have to show $a\in SSC(b)$; but that follows from $a\in SSC(a)$ and the definition of subset. That completes the right to left direction. That completes the proof of the lemma.
@@
Let $b$ be a finite set. Then the subset relation on $SSC(b)$ is decidable. That is,
$\forall x,y\in SSC(b)\,(x\subseteq y\ \lor\ x\not\subseteq y).$
Proof. Assume $b\in{\sf FINITE}$. By Lemma 3.17, $SSC(b)\in{\sf FINITE}$. Then by Lemma 3.3,
(54) $SSC(b)\in{\sf DECIDABLE}$
@@
(55) $y\in SSC(b)\ {\rightarrow}\ \forall x\in SSC(b)\,(x\subseteq y\ \lor\ x\not\subseteq y).$
@@
Base case. When $y=\Lambda$, we will prove
$\forall x\in SSC(b)\,(x\subseteq\Lambda\lor x\not\subseteq\Lambda).$
Assume $x\in SSC(b)$. We have $x\subseteq\Lambda$ if and only if $x=\Lambda$, so it suffices to prove $x=\Lambda\ \lor\ x\neq\Lambda$. But that follows from (54). That completes the base case.
Induction step. Let $y=z\cup\{c\}$, with $c\not\in z$ and $z\in SSC(b)$ and $y\subseteq b$. Then $c\in b$ The induction hypothesis is
(56) $z\in SSC(b)\ {\rightarrow}\ \forall x\in SSC(b)\,(x\subseteq z\ \lor\ x\not\subseteq z).$
@@
(57) $y\in SSC(b)\ {\rightarrow}\ \forall x\in SSC(b)\,(x\subseteq y\ \lor\ x\not\subseteq y)$
Assume $y\in SSC(b)$ and $x\in SSC(b)$. We have to prove $x\subseteq y\ \lor\ x\not\subseteq y$. That is,
@@
$y\in SSC(b)$   assumed above
$z\cup\{c\}\in SSC(b)$   since $y=z\cup\{c\}$
I say that $z\in SSC(b)$. To prove that, let $u\in z$. Since $z\cup\{c\}\in SSC(b)$, $u\in z\cup\{c\}\ \lor\ u\not\in z\cup\{c\}$. Since $c\not\in z$, $u\neq c$. Therefore $u\in z\lor u\not\in z$. Then $z\in SSC(b)$ as claimed.
I say that also $x-\{c\}\in SSC(b)$. Since $b$ is finite, it has decidable equality by Lemma 3.3. Then for $y\in b$, we have $y=c\ \lor\ y\neq c$. Since $x\in SSC(b)$ we have $y\in x\ \lor\ y\not\in x$. Then a short argument by cases shows $y\in x-\{c\}\ \lor\ y\not\in x\{c\}$. Then $x-\{c\}\in SSC(b)$, as claimed.
By (56) and $z\in SSC(b)$, we have
(58) $\forall x\in SSC(b)\,(x\subseteq z\ \lor\ x\not\subseteq z).$
Since $x\in SSC(b)$, we have $c\in x\ \lor\ c\not\in x$. We argue by cases accordingly.
Case 1: $c\in x$. Then $x\subseteq z\cup\{c\}$ if and only if $x-\{c\}\subseteq z$. By (58), instantiated to $x-\{c\}$ in place of $x$ (which is allowed since $x-\{c\}\in SSC(b)$), we have
@@
Case 2: $c\not\in x$. Then $x\subseteq z\cup\{c\}{\ \leftrightarrow\ }x\subseteq z$, so (57) follows from the induction hypothesis (56). That completes Case 1. That completes the induction step. That completes the proof of the lemma.
@@
$a\in SSC(b)\ {\rightarrow}\ SSC(a)\in SSC(SSC(b)).$
Proof. Suppose $a\in SSC(b)$. By Lemma 6.12,
(59) $SSC(a)\subseteq SSC(b)$
It remains to show that $SSC(a)$ is a separable subset of $SSC(b)$; that is,
$SSC(b)=SSC(a)\cup(SSC(b)-SSC(a)).$
@@
[…]
@@
Ad the left-to-right direction of (60): suppose $t\in SSC(b)$. Then $t\subseteq b$. By Lemma 6.13,
(61) $t\subseteq a\lor t\not\subseteq a.$
It remains to prove the left-to-right direction of (60). Suppose $t\in SSC(b)$. We argue by cases using (61).
Case 1: $t\subseteq a$. It suffices to prove $t\in SSC(a)$. It remains to prove $a=t\cup(a-t)$. We have
$\forall u\in b\,(u\in t\ \lor\ u\not\in t)\mbox{\qquad since $t\in SSC(b)$}$
@@
Case 2: $t\not\subseteq a$. Then $t\not\in SSC(a)$. Since $t\in SSC(b)$, the second disjunct on the right of (60) holds. That completes Case 2. That completes the proof of the lemma.
@@
$USC(a\cup\{c\})=USC(a)\cup\{\{c\}\}.$
Proof. By extensionality it suffices to verify the two sides have the same members.
Left to right: Let $x\in USC(a\cup\{c\})$. Then $x=\{u\}$ for some $u\in a\cup\{c\}$. Then $u\in a\ \lor\ a=c$. If $u\in a$ then $x\in USC(a)$ and hence $x\in USC(a)\cup\{\{c\}\}$. That completes the left-to-right direction.
Right to left: Let $x\in USC(a)\cup\{\{c\}\}$. Then $x\in USC(a)\ \lor\ x=\{c\}$. If $x\in USC(a)$, then $x\in USC(a\cup\{c\})$ by Lemma 6.10. If $x=\{c\}$, then $x\in USC(a\cup\{c\})$ by definition of USC. That completes the proof of the lemma.
@@
$USC(a-b)=USC(a)-USC(b).$
Proof. By the definitions of USC and set difference, using about 50 straightforward inferences, which we choose to omit.
@@
$USC(\Lambda)=\Lambda$.
Proof. Suppose $x\in USC(\Lambda)$. By definition of $USC$, there exists $a\in\Lambda$ such that $x=\{a\}$. But that contradicts the definition of $\Lambda$. That completes the proof of the lemma.
@@
$x\in a{\ \leftrightarrow\ }\{x\}\in USC(a).$
Proof. Left to right, by definition of $USC(a)$. Right to left: if $\{x\}\in USC(a)$, then for some $y\in a$, $\{x\}=\{y\}$. Then by extensionality $x=y$. That completes the proof of the lemma.
@@
$SSC(\Lambda)=\{\Lambda\}$.
Proof. The only subset of $\Lambda$ is $\Lambda$, and it is a stable subset.
@@
Proof. Let $f:a\to b$ be a similarity. Since $a$ is inhabited, there exists some $c\in a$. Fix $c$. Then $f(c)\in b$. Hence $b$ is inhabited. That completes the proof of the lemma.
@@
Proof. The left-to-right direction is logically valid. We prove the right-to-left implication by induction on $y$. The formula to be proved is stratified, giving $x$ and $y$ index 0, so induction is legal.
@@
$\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ \lor\ x=y)\ {\rightarrow}\ \neg\neg\,P(x)$   by 62)
$\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ {\rightarrow}\ \neg\neg\,x\in P\ \land\ x=y\ {\rightarrow}\ \neg\neg\,P(x))$   by logic
$\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ {\rightarrow}\ \neg\neg\,x\in P)\ \land\ \neg\neg\,y\in P$   by logic
$\neg\neg\,\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ {\rightarrow}\ x\in P)\ \land\ \neg\neg\,y\in P$   induction hyp.
$\neg\neg\,\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ x\leq y\ {\rightarrow}\ x\in P)$   by 62)
That completes the induction step. That completes the proof of the lemma.
7. Cardinal exponentiation
Specker 4.1 follows Rosser in defining $2^{m}$ for cardinals $m$. They define $2^{m}$ to be the cardinal of $SC(a)$ where $USC(a)\in m$. That definition requires some modification to be of use constructively. It is separable subsets of $a$ that correspond to functions from $a$ to 2, so it makes sense to use $SSC(a)$, the class of separable subsets of $a$, instead of $SC(a)$.
@@ 7. Cardinal exponentiation
$2^{m}=\{u:\exists a\,(USC(a)\in m\ \land\ u\sim SSC(a)\}.$
@@
Let $m\in{\mathbb{F}}$ and $USC(a)\in m$. Then $SSC(a)\in 2^{m}$, and $2^{m}=Nc(SSC(a))$.
Remark. This is Specker’s definition of $2^{m}$, but our definition avoids a case distinction as to whether $m$ does or does not contain a set of the form $USC(a)$.
Proof. Suppose $m\in{\mathbb{F}}$ and $USC(a)\in m$. I say that $2^{m}$ is a cardinal, i.e., it is closed under similarity. Suppose $u$ and $v$ are members of $2^{m}$. Then there exist $a$ and $b$ such that $USC(a)$ and $USC(b)$ are both in $m$ and $u\sim SSC(a)$ and $v\sim SSC(b)$. Then by Lemma 4.9, $USC(a)\sim USC(b)$. By Lemma 6.7, $a\sim b$. By Lemma 6.8, $SSC(a)\sim SSC(b)$. By Lemma 2.11, $u\sim v$. Hence, as claimed, $2^{m}$ is a cardinal.
Therefore $2^{m}$ and $Nc(SSC(a))$ are both closed under similarity. Since they both contain $SSC(a)$, they each consist of all sets similar to $SSC(a)$. Hence by extensionality, they are equal. That completes the proof of the lemma.
Remark. We note that $2^{m}\neq\Lambda$ does not a priori imply that $2^{m}$ is inhabited, so we must carefully distinguish these two statements as hypotheses of lemmas. $2^{m}$ is inhabited if $m$ contains a set of the form $USC(a)$. $2^{m}\neq\Lambda$ means not-not $m$ contains such a set.
@@
is definable in INF.
Proof. We have to show that the relation is definable by a formula that can be stratified, giving the two members of ordered pairs the same index. The formula in Definition 7.1 is
$2^{m}=\{u:\exists a\,(USC(a)\in m\ \land\ u\sim SSC(a)\}.$
Stratify it, giving $a$ index 0, $USC(a)$ and $SSC(a)$ and $u$ index 1, $m$ index 2. Then $2^{m}$ gets one index higher than $u$, namely 2, which is the same index that $m$ gets. That completes the proof.
@@
If $m\in{\mathbb{F}}$ and $2^{m}$ is inhabited, then there exists $a$ such that $USC(a)\in m$ and $SSC(a)\in 2^{m}$.
Proof. Suppose $2^{m}$ is inhabited. Then by Definition 7.1, there exists $a$ with $USC(a)\in m$, and $2^{m}$ contains any set similar to $SSC(a)$. Since $SSC(a)\sim SSC(a)$, by Lemma 2.11, we have $SSC(a)\in 2^{m}$. That completes the proof of the lemma.
@@
Proof. Suppose $m$ is a finite cardinal and $2^{m}$ is inhabited. By Definition 7.1, there exists $a$ such that $USC(a)\in m$ and $SSC(a)\in 2^{m}$. Then $Nc(SSC(a))=2^{m}$, by Definition 7.1. We have
$USC(a)\in{\sf FINITE}$   by Lemma 4.4
@@
$SSC(a)\in{\sf FINITE}$   by Lemma 3.17
$Nc(SSC(a))\in{\mathbb{F}}$   by Lemma 4.21
$2^{m}\in{\mathbb{F}}$   since $Nc(SSC(a))=2^{m}$
That completes the proof of the lemma.
@@
Proof. ${\sf zero}=\{\Lambda\}$. It therefore contains $\Lambda=USC(\Lambda)$. Hence $2^{\sf zero}$ is inhabited and contains $SSC(\Lambda)$. But $\Lambda$ has only one subset, namely $\Lambda$, which is a separable subset, so $SSC(\Lambda)=\{\Lambda\}={\sf zero}$. Thus $2^{\sf zero}=Nc({\sf zero})={\sf one}$. That completes the proof of the lemma.
@@
Proof. ${\sf one}$ is the set of all singletons. It therefore contains $\{{\sf zero}\}=USC({\sf zero})$. Then $2^{\sf one}$ contains $SSC({\sf zero})$. There are exactly two subsets of $\{\Lambda\}$, namely $\Lambda$ and $\{\Lambda\}$, and both are separable. Hence $2^{\sf one}$ contains the two-element set $SSC({\sf zero})=\{\Lambda,\{\Lambda\}\}$. That set belongs to ${\sf two}={\sf one}^{+}$ since it is equal to $\{\{\Lambda\}\}\cup\{\Lambda\}$, and the singleton $\{\{\Lambda\}\}$ belongs to ${\sf one}$ and $\{\Lambda\}\not\in\{\{\Lambda\}\}$. Therefore $2^{\sf one}$ and ${\sf one}^{+}$ have a common element. Both are cardinals, by Lemma 7.5. Then by Lemma 5.7, $2^{\sf one}={\sf two}$. That completes the proof of the lemma.
@@
Proof. By definition, ${\sf four}={\sf three}^{+}={\sf two}^{++}$. One can show (but we omit the details) that
$USC(\{{\sf one},{\sf two}\})=\{\{{\sf one}\},\{{\sf two}\}\}\in{\sf two}$

### v6

@@ 6 Power sets and similarity
$t\in\mathcal{P}_{1}(b)$ ${\ \leftrightarrow\ }$ $t\in\mathcal{P}_{1}(a)\ \lor\ (t\in\mathcal{P}_{1}(b)\ \land\ t\not\in\mathcal{P}_{1}(a)).$ (53)
Then we need only consider unit classes $t=\{x\}$, and using the fact that $\{x\}\in\mathcal{P}_{1}(b){\ \leftrightarrow\ }t\in b$, and $\{x\}\in\mathcal{P}_{1}(a){\ \leftrightarrow\ }t\in a$, (53) follows from (51). ∎
@@
$a\in\mathcal{P}_{s}(b){\ \leftrightarrow\ }\mathcal{P}_{s}(a)\subseteq\mathcal{P}_{s}(b).$
Proof.
Left to right: Suppose $a\in\mathcal{P}_{s}(b)$. Then $a\subseteq b$ and
$b$ $=$ $a\cup(b-a)$ (54)
Now let $x\in\mathcal{P}_{s}(a)$. We must show $x\in\mathcal{P}_{s}(b)$. Since $x\in\mathcal{P}_{s}(a)$, we have $x\subseteq a$. Since $a\subseteq b$ we have $x\subseteq b$. We have
$x\in\mathcal{P}_{s}(a)$
$a=x\cup(a-x)$   by definition of $\mathcal{P}_{s}(a)$
@@
Right to left: Suppose $\mathcal{P}_{s}(a)\subseteq\mathcal{P}_{s}(b)$. We have to show $a\in\mathcal{P}_{s}(b)$; but that follows from $a\in\mathcal{P}_{s}(a)$ and the definition of subset. That completes the right to left direction. ∎
@@
Let $b$ be a finite set. Then the subset relation on $\mathcal{P}_{s}(b)$ is decidable. That is,
$\forall x,y\in\mathcal{P}_{s}(b)\,(x\subseteq y\ \lor\ x\not\subseteq y).$
Proof.
Assume $b\in{\sf FINITE}$. By Lemma 3.17, $\mathcal{P}_{s}(b)\in{\sf FINITE}$. Then by Lemma 3.3,
$\mathcal{P}_{s}(b)\in{\sf DECIDABLE}$ (55)
@@
$y\in\mathcal{P}_{s}(b)\ {\rightarrow}\ \forall x\in\mathcal{P}_{s}(b)\,(x\subseteq y\ \lor\ x\not\subseteq y).$ (56)
@@
Base case. When $y=\emptyset$, we will prove
$\forall x\in\mathcal{P}_{s}(b)\,(x\subseteq\emptyset\lor x\not\subseteq\emptyset).$
Assume $x\in\mathcal{P}_{s}(b)$. We have $x\subseteq\emptyset$ if and only if $x=\emptyset$, so it suffices to prove $x=\emptyset\ \lor\ x\neq\emptyset$. But that follows from (55). That completes the base case.
Induction step. Let $y=z\cup\{c\}$, with $c\not\in z$ and $z\in\mathcal{P}_{s}(b)$ and $y\subseteq b$. Then $c\in b$ The induction hypothesis is
$z\in\mathcal{P}_{s}(b)\ {\rightarrow}\ \forall x\in\mathcal{P}_{s}(b)\,(x\subseteq z\ \lor\ x\not\subseteq z).$ (57)
@@
$y\in\mathcal{P}_{s}(b)\ {\rightarrow}\ \forall x\in\mathcal{P}_{s}(b)\,(x\subseteq y\ \lor\ x\not\subseteq y)$ (58)
Assume $y\in\mathcal{P}_{s}(b)$ and $x\in\mathcal{P}_{s}(b)$. We have to prove $x\subseteq y\ \lor\ x\not\subseteq y$. That is,
@@
$y\in\mathcal{P}_{s}(b)$   assumed above
$z\cup\{c\}\in\mathcal{P}_{s}(b)$   since $y=z\cup\{c\}$
I say that $z\in\mathcal{P}_{s}(b)$. To prove that, let $u\in z$. Since $z\cup\{c\}\in\mathcal{P}_{s}(b)$, $u\in z\cup\{c\}\ \lor\ u\not\in z\cup\{c\}$. Since $c\not\in z$, $u\neq c$. Therefore $u\in z\lor u\not\in z$. Then $z\in\mathcal{P}_{s}(b)$ as claimed.
I say that also $x-\{c\}\in\mathcal{P}_{s}(b)$. Since $b$ is finite, it has decidable equality by Lemma 3.3. Then for $y\in b$, we have $y=c\ \lor\ y\neq c$. Since $x\in\mathcal{P}_{s}(b)$ we have $y\in x\ \lor\ y\not\in x$. Then a short argument by cases shows $y\in x-\{c\}\ \lor\ y\not\in x\{c\}$. Then $x-\{c\}\in\mathcal{P}_{s}(b)$, as claimed.
By (57) and $z\in\mathcal{P}_{s}(b)$, we have
$\forall x\in\mathcal{P}_{s}(b)\,(x\subseteq z\ \lor\ x\not\subseteq z).$ (59)
Since $x\in\mathcal{P}_{s}(b)$, we have $c\in x\ \lor\ c\not\in x$. We argue by cases accordingly.
Case 1: $c\in x$. Then $x\subseteq z\cup\{c\}$ if and only if $x-\{c\}\subseteq z$. By (59), instantiated to $x-\{c\}$ in place of $x$ (which is allowed since $x-\{c\}\in\mathcal{P}_{s}(b)$), we have
@@
Case 2: $c\not\in x$. Then $x\subseteq z\cup\{c\}{\ \leftrightarrow\ }x\subseteq z$, so (58) follows from the induction hypothesis (57). That completes Case 1. That completes the induction step. ∎
@@
$a\in\mathcal{P}_{s}(b)\ {\rightarrow}\ \mathcal{P}_{s}(a)\in\mathcal{P}_{s}(\mathcal{P}_{s}(b)).$
Proof.
Suppose $a\in\mathcal{P}_{s}(b)$. By Lemma 6.12,
$\mathcal{P}_{s}(a)\subseteq\mathcal{P}_{s}(b)$ (60)
It remains to show that $\mathcal{P}_{s}(a)$ is a separable subset of $\mathcal{P}_{s}(b)$; that is,
$\mathcal{P}_{s}(b)=\mathcal{P}_{s}(a)\cup(\mathcal{P}_{s}(b)-\mathcal{P}_{s}(a)).$
@@
$t\in\mathcal{P}_{s}(b)$ ${\ \leftrightarrow\ }$ $t\in\mathcal{P}_{s}(a)\ \lor\ (t\in\mathcal{P}_{s}(b)\ \land\ t\not\in\mathcal{P}_{s}(a))$ (61)
@@
Ad the left-to-right direction of (61): suppose $t\in\mathcal{P}_{s}(b)$. Then $t\subseteq b$. By Lemma 6.13,
$t\subseteq a\lor t\not\subseteq a.$ (62)
It remains to prove the left-to-right direction of (61). Suppose $t\in\mathcal{P}_{s}(b)$. We argue by cases using (62).
Case 1: $t\subseteq a$. It suffices to prove $t\in\mathcal{P}_{s}(a)$. It remains to prove $a=t\cup(a-t)$. We have
$\forall u\in b\,(u\in t\ \lor\ u\not\in t)\mbox{\qquad since $t\in\mathcal{P}_{s}(b)$}$
@@
Case 2: $t\not\subseteq a$. Then $t\not\in\mathcal{P}_{s}(a)$. Since $t\in\mathcal{P}_{s}(b)$, the second disjunct on the right of (61) holds. That completes Case 2. ∎
@@
$\mathcal{P}_{1}(a\cup\{c\})=\mathcal{P}_{1}(a)\cup\{\{c\}\}.$
Proof.
By extensionality it suffices to verify the two sides have the same members.
Left to right: Let $x\in\mathcal{P}_{1}(a\cup\{c\})$. Then $x=\{u\}$ for some $u\in a\cup\{c\}$. Then $u\in a\ \lor\ a=c$. If $u\in a$ then $x\in\mathcal{P}_{1}(a)$ and hence $x\in\mathcal{P}_{1}(a)\cup\{\{c\}\}$. That completes the left-to-right direction.
Right to left: Let $x\in\mathcal{P}_{1}(a)\cup\{\{c\}\}$. Then $x\in\mathcal{P}_{1}(a)\ \lor\ x=\{c\}$. If $x\in\mathcal{P}_{1}(a)$, then $x\in\mathcal{P}_{1}(a\cup\{c\})$ by Lemma 6.10. If $x=\{c\}$, then $x\in\mathcal{P}_{1}(a\cup\{c\})$ by definition of $\mathcal{P}_{1}$. ∎
@@
$\mathcal{P}_{1}(a-b)=\mathcal{P}_{1}(a)-\mathcal{P}_{1}(b).$
Proof.
By the definitions of $\mathcal{P}_{1}$ and set difference, using about 50 straightforward inferences, which we choose to omit. ∎
@@
$\mathcal{P}_{1}(\emptyset)=\emptyset$.
Proof.
Suppose $x\in\mathcal{P}_{1}(\emptyset)$. By definition of $\mathcal{P}_{1}$, there exists $a\in\emptyset$ such that $x=\{a\}$. But that contradicts the definition of $\emptyset$. ∎
@@
$x\in a{\ \leftrightarrow\ }\{x\}\in\mathcal{P}_{1}(a).$
Proof.
Left to right, by definition of $\mathcal{P}_{1}(a)$. Right to left: if $\{x\}\in\mathcal{P}_{1}(a)$, then for some $y\in a$, $\{x\}=\{y\}$. Then by extensionality $x=y$. ∎
@@
$\mathcal{P}_{s}(\emptyset)=\{\emptyset\}$.
Proof.
The only subset of $\emptyset$ is $\emptyset$, and it is a separable subset. ∎
@@
Proof.
Let $f:a\to b$ be a similarity. Since $a$ is inhabited, there exists some $c\in a$. Fix $c$. Then $f(c)\in b$. Hence $b$ is inhabited. ∎
@@
Proof.
The left-to-right direction is logically valid. We prove the right-to-left implication by induction on $y$. The formula to be proved is stratified, giving $x$ and $y$ index 0, so induction is legal.
@@
$\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ \lor\ x=y)\ {\rightarrow}\ \neg\neg\,x\in P)$   by (63)
$\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ {\rightarrow}\ \neg\neg\,x\in P)\ \land\ (x=y\ {\rightarrow}\ \neg\neg\,x\in P))$   by logic
$\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ {\rightarrow}\ \neg\neg\,x\in P))\ \land\ \neg\neg\,y\in P$   by logic
$\neg\neg\,\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ (x<y\ {\rightarrow}\ x\in P))\ \land\ \neg\neg\,y\in P$   induction hyp.
$\neg\neg\,\forall x\,(x\in{\mathbb{F}}\ {\rightarrow}\ x\leq y\ {\rightarrow}\ x\in P)$   by (63)
That completes the induction step. ∎
7 Cardinal exponentiation
Specker 4.1 follows Rosser in defining $2^{m}$ for cardinals $m$. They define $2^{m}$ to be the cardinal of $\mathcal{P}(a)$ where $\mathcal{P}_{1}(a)\in m$. That definition requires some modification to be of use constructively. It is separable subsets of $a$ that correspond to functions from $a$ to 2, so it makes sense to use $\mathcal{P}_{s}(a)$, the class of separable subsets of $a$, instead of $\mathcal{P}(a)$.
@@ 7 Cardinal exponentiation
$2^{m}=\{u:\exists a\,(\mathcal{P}_{1}(a)\in m\ \land\ u\sim\mathcal{P}_{s}(a))\}.$
@@
Let $m\in{\mathbb{F}}$ and $\mathcal{P}_{1}(a)\in m$. Then $\mathcal{P}_{s}(a)\in 2^{m}$, and $2^{m}=|\mathcal{P}_{s}(a|)$.
Remark. This is Specker’s definition of $2^{m}$, but our definition avoids a case distinction as to whether $m$ does or does not contain a set of the form $\mathcal{P}_{1}(a)$.
Proof.
Suppose $m\in{\mathbb{F}}$ and $\mathcal{P}_{1}(a)\in m$. I say that $2^{m}$ is a cardinal, i.e., it is closed under similarity. Suppose $u$ and $v$ are members of $2^{m}$. Then there exist $a$ and $b$ such that $\mathcal{P}_{1}(a)$ and $\mathcal{P}_{1}(b)$ are both in $m$ and $u\sim\mathcal{P}_{s}(a)$ and $v\sim\mathcal{P}_{s}(b)$. Then by Lemma 4.9, $\mathcal{P}_{1}(a)\sim\mathcal{P}_{1}(b)$. By Lemma 6.7, $a\sim b$. By Lemma 6.8, $\mathcal{P}_{s}(a)\sim\mathcal{P}_{s}(b)$. By Lemma 2.11, $u\sim v$. Hence, as claimed, $2^{m}$ is a cardinal.
Therefore $2^{m}$ and $|\mathcal{P}_{s}(a|)$ are both closed under similarity. Since they both contain $\mathcal{P}_{s}(a)$, they each consist of all sets similar to $\mathcal{P}_{s}(a)$. Hence by extensionality, they are equal. ∎
Remark. We note that $2^{m}\neq\emptyset$ does not a priori imply that $2^{m}$ is inhabited, so we must carefully distinguish these two statements as hypotheses of lemmas. $2^{m}$ is inhabited if $m$ contains a set of the form $\mathcal{P}_{1}(a)$. $2^{m}\neq\emptyset$ means not-not $m$ contains such a set.
Discussion. It is possible, of course, to investigate what happens if we use intuitionistic logic, but keep the classical definitions of order and exponentiation. The most obvious difficulty with this approach is that the integers $\mathbb{F}$} are not closed under exponentiation. For example, let us calculate what $2^{{\sf one}}$ would be. We have $\{\{\emptyset\}\}=\mathcal{P}_{1}(\{\emptyset\})\in{\sf one}$. So $2^{{\sf zero}}$ would be the cardinal of $SC(\{\emptyset)\}$, instead of the cardinal of $\mathcal{P}_{s}(\{\emptyset\})$. But $SC(\{\emptyset\})$ contains every set of the form $X_{P}=\{x:x=\emptyset\ \land\ P\}$, where $P$ is a stratified formula not containing the variable $x$. Unless we can prove or refute $P$, we cannot prove that $X_{P}$ is one of the two members of $\mathcal{P}_{s}(\emptyset)$, and in fact there is no hope of proving $2^{{\sf one}}$ is an integer. Hence this notion is useless for constructive mathematics in NF. Still we did investigate the matter further, as there was the possibility that this approach might help analyze Specker’s proof. In short, it did not help. Without the axiom of choice, one can prove nothing useful about large cardinals. For example, one cannot prove $2^{x}=2^{y}\ {\rightarrow}\ x=y$ for cardinals; there might even be incomparable $x,y$ such that $2^{x}=2^{y}$. That might even be the case with $2^{x}=2^{\kappa}=\kappa$, where $\kappa$ is the cardinal of $\mathbb{V}$}. We consider this subject no further.
@@
is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
Proof.
We have to show that the relation is definable by a formula that can be stratified, giving the two members of ordered pairs the same index. The formula in Definition 7.1 is
$2^{m}=\{u:\exists a\,(\mathcal{P}_{1}(a)\in m\ \land\ u\sim\mathcal{P}_{s}(a)\}.$
Stratify it, giving $a$ index 0, $\mathcal{P}_{1}(a)$ and $\mathcal{P}_{s}(a)$ and $u$ index 1, $m$ index 2. Then $2^{m}$ gets one index higher than $u$, namely 2, which is the same index that $m$ gets. ∎
@@
If $2^{m}$ is inhabited, then there exists $a$ such that $\mathcal{P}_{1}(a)\in m$ and $\mathcal{P}_{s}(a)\in 2^{m}$.
Proof.
Suppose $2^{m}$ is inhabited. Then by Definition 7.1, there exists $a$ with $\mathcal{P}_{1}(a)\in m$, and $2^{m}$ contains any set similar to $\mathcal{P}_{s}(a)$. Since $\mathcal{P}_{s}(a)\sim\mathcal{P}_{s}(a)$, by Lemma 2.11, we have $\mathcal{P}_{s}(a)\in 2^{m}$. ∎
@@
Proof.
Suppose $m$ is a finite cardinal and $2^{m}$ is inhabited. By Definition 7.1, there exists $a$ such that $\mathcal{P}_{1}(a)\in m$ and $\mathcal{P}_{s}(a)\in 2^{m}$. Then $|\mathcal{P}_{s}(a|)=2^{m}$, by Definition 7.1. We have
$\mathcal{P}_{1}(a)\in{\sf FINITE}$   by Lemma 4.4
@@
$\mathcal{P}_{s}(a)\in{\sf FINITE}$   by Lemma 3.17
$|\mathcal{P}_{s}(a|)\in{\mathbb{F}}$   by Lemma 4.21
$2^{m}\in{\mathbb{F}}$   since $|\mathcal{P}_{s}(a|)=2^{m}$
@@
Proof.
$\sf zero$$\{\emptyset\}$. It therefore contains $\emptyset=\mathcal{P}_{1}(\emptyset)$. Hence $2^{\sf zero}$ is inhabited and contains $\mathcal{P}_{s}(\emptyset)$. But $\emptyset$ has only one subset, namely $\emptyset$, which is a separable subset, so $\mathcal{P}_{s}(\emptyset)=\{\emptyset\}={\sf zero}$. Thus $2^{\sf zero}=|{\sf zero}|={\sf one}$. ∎
@@
Proof.
$\sf one$} is the set of all singletons. It therefore contains $\{{\sf zero}\}=\mathcal{P}_{1}({\sf zero})$. Then $2^{\sf one}$ contains $\mathcal{P}_{s}({\sf zero})$. There are exactly two subsets of $\{\emptyset\}$, namely $\emptyset$ and $\{\emptyset\}$, and both are separable. Hence $2^{\sf one}$ contains the two-element set $\mathcal{P}_{s}({\sf zero})=\{\emptyset,\{\emptyset\}\}$. That set belongs to ${\sf two}={\sf one}^{+}$ since it is equal to $\{\{\emptyset\}\}\cup\{\emptyset\}$, and the singleton $\{\{\emptyset\}\}$ belongs to $\sf one$} and $\{\emptyset\}\not\in\{\{\emptyset\}\}$. Therefore $2^{\sf one}$ and $\sf one$$^{+}$ have a common element. Both are cardinals, by Lemma 7.5. Then by Lemma 4.24, $2^{\sf one}={\sf two}$. ∎
@@
Proof.
By definition, ${\sf four}={\sf three}^{+}={\sf two}^{++}$. One can show (but we omit the details) that
$\mathcal{P}_{1}(\{{\sf one},{\sf two}\})=\{\{{\sf one}\},\{{\sf two}\}\}\in{\sf two}$